The sales, income from operations, and invested assets for each division of Jackson Corporation are as follows:

Sales

Income from

Operations

Invested

Assets

Division E

$4,000,000

$550,000

$2,400,000

Division F

4,800,000

760,000

2,500,000

Division G

7,000,000

860,000

2,800,000

Using the expanded expression, determine the profit margin, investment turnover, and rate of return on investment for each division. Round answers to one decimal place.

Which Division is the most profitable per dollar invested?
Answer #1
Formula Division E Division F Division G
Profit margin Operating income/ Sales $550000*100/4000000= 13.75% $760000*100/4800000= 15.83% $860000*100/7000000= 12.29%
Investment turnover Sales/ Invested assets $4000000/2400000= 1.67 times $4800000/2500000= 1.92 times $7000000/2800000= 2.5 times
Rate of return on investment Operating income/ Invested assets $550000*100/2400000= 22.92% $760000*100/2500000= 30.4% $860000*100/2800000= 30.71%

Which Division is the most profitable per dollar invested?

Division G as division G has the highest Rate of return on Investment